# Web Artifact Attacks Disrupt Vision Language Models

### v1

@@ Appendix B Attack Sucess Rate by Location
In Sec. 3.1.3, we discussed how we optimize the artifact location placement as part of our attack. Fig. 11 examines how the placement of artifacts affects attack success rates across different artifact types. The results indicate that artifacts positioned in the top-center region of the image consistently lead to higher misclassification rates, with text-based artifacts being the most effective. This suggests that vision-language models prioritize information in certain spatial regions, likely due to biases in pretraining datasets, where text frequently appears near the top of images (e.g., headlines, labels, or watermarks). Graphics with embedded text also show higher success rates in the top-center, though to a lesser extent than pure text, while graphics without text have a more uniform but overall lower effect. These findings highlight the importance of spatial biases in model vulnerability and suggest that adversarial manipulations may be optimized further by strategically placing artifacts in regions that models inherently attend to more strongly.
@@ Appendix C Attack Success by Dataset
In Sec. 4.1, we report the average performance over 5 datasets. In this Section, we break down performance by dataset. Fig. 12 shows that artifact-based attacks achieve higher success rates in datasets containing human-related attributes (FairFace [14], CelebA [19]) compared to non-human classification tasks (Aircraft [20], Country211 [26]). This trend is particularly pronounced for text-based artifacts, which consistently lead to the highest misclassification rates in human-related datasets. One possible explanation is that vision-language models are pretrained on web-scale data where text often co-occurs with human-related concepts, reinforcing spurious associations between textual artifacts and human characteristics. In contrast, non-human tasks like aircraft recognition rely more on fine-grained visual details, making them less susceptible to artifacts that exploit text or graphical elements. These findings suggest that artifact-based attacks pose a greater risk to applications involving human-centric classifications, where models may rely more heavily on text-based biases.
Appendix D Effect of Transparency and Size on Attack Success Rate
In Section 4.1., we fix the artifact size to 10th of the image sizs and transparency to 1.0. In this Section, we ablate both settings. Fig. 13 evaluates how reducing artifact size and transparency affects attack success rates. The results show that as artifacts become smaller or more transparent, their effectiveness declines across all artifact types, with text-based artifacts remaining the most resilient. This suggests that larger and more visible artifacts are more likely to be leveraged by the model, while smaller or faded artifacts are either ignored or contribute less to misclassification. Notably, the steepest decline occurs for graphics-based artifacts, particularly those without text, indicating that purely visual artifacts are more sensitive to reductions in size and visibility. These findings highlight that while reducing artifact saliency can mitigate their impact, text-based artifacts still pose a considerable threat, even when minimally visible.
@@ Appendix E Descriptions of Artifacts for Mitigation
Fig. 14 illustrates how automatically generated captions are incorporated into our artifact-aware prompting strategy (see Sec 5) to mitigate artifact-based attacks. Inspired by prior work [7], which demonstrated that Vision-Language Models (VLMs) can adjust their attention when given more informative prompts, we extend this approach to graphical artifacts. Unlike text-based artifacts, which can be directly embedded into prompts, graphical artifacts lack an explicit textual representation, making them harder for the model to explicitly consider. To address this, we generate structured descriptions of graphical symbols and append them to the input prompt, ensuring that the model processes them explicitly rather than forming unintended associations.
The captions in Fig. 14 serve this purpose by neutralizing spurious correlations and guiding the model toward actual visual semantics. For example, instead of allowing the model to infer associations based on dataset biases (e.g., linking a hazard symbol to danger-related concepts), the caption describes it as “a yellow sign with a black hazard symbol”, removing any loaded interpretation.

### v2

@@ Appendix B Attack Sucess Rate by Location
In Sec. 3.1.3, we discussed how we optimize the artifact location placement as part of our attack. Fig.˜12 examines how the placement of artifacts affects attack success rates across different artifact types. The results indicate that artifacts positioned in the top-center region of the image consistently lead to higher misclassification rates, with text-based artifacts being the most effective. This suggests that vision-language models prioritize information in certain spatial regions, likely due to biases in pretraining datasets, where text frequently appears near the top of images (e.g., headlines, labels, or watermarks). Graphics with embedded text also show higher success rates in the top-center, though to a lesser extent than pure text, while graphics without text have a more uniform but overall lower effect. These findings highlight the importance of spatial biases in model vulnerability and suggest that adversarial manipulations may be optimized further by strategically placing artifacts in regions that models inherently attend to more strongly.
@@ Appendix C Attack Success by Dataset
In Sec. 4.1, we report the average performance over 5 datasets. In this Section, we break down performance by dataset. Fig.˜13 shows that Web Artifact Attacks achieve higher success rates in datasets containing human-related attributes (FairFace [15], CelebA [20]) compared to non-human classification tasks (Aircraft [21], Country211 [28]). This trend is particularly pronounced for text-based artifacts, which consistently lead to the highest misclassification rates in human-related datasets. One possible explanation is that vision-language models are pretrained on web-scale data where text often co-occurs with human-related concepts, reinforcing spurious associations between textual artifacts and human characteristics. In contrast, non-human tasks like aircraft recognition rely more on fine-grained visual details, making them less susceptible to artifacts that exploit text or graphical elements. These findings suggest that Web Artifact Attacks pose a greater risk to applications involving human-centric classifications, where models may rely more heavily on text-based biases.
Appendix D Effect of Artifact’s Transparency and Size
In Sec. 4.1., we fix the artifact size to 10th of the image sizs and transparency to 1.0. In this Section, we ablate both settings. Fig.˜14 evaluates how reducing artifact size and transparency affects attack success rates. The results show that as artifacts become smaller or more transparent, their effectiveness declines across all artifact types, with text-based artifacts remaining the most resilient. This suggests that larger and more visible artifacts are more likely to be leveraged by the model, while smaller or faded artifacts are either ignored or contribute less to misclassification. Notably, the steepest decline occurs for graphics-based artifacts, particularly those without text, indicating that purely visual artifacts are more sensitive to reductions in size and visibility. These findings highlight that while reducing artifact saliency can mitigate their impact, text-based artifacts still pose a considerable threat, even when minimally visible.
@@ Appendix E Descriptions of Artifacts for Mitigation
Fig.˜15 illustrates how automatically generated captions are incorporated into our artifact-aware prompting strategy (see Sec 5) to mitigate Web Artifact Attacks. Inspired by prior work [8], which demonstrated that Vision-Language Models (VLMs) can adjust their attention when given more informative prompts, we extend this approach to graphical artifacts. Unlike text-based artifacts, which can be directly embedded into prompts, graphical artifacts lack an explicit textual representation, making them harder for the model to explicitly consider. To address this, we generate structured descriptions of graphical symbols and append them to the input prompt, ensuring that the model processes them explicitly rather than forming unintended associations.
The captions in Fig.˜15 serve this purpose by neutralizing spurious correlations and guiding the model toward actual visual semantics. For example, instead of allowing the model to infer associations based on dataset biases (e.g., linking a hazard symbol to danger-related concepts), the caption describes it as “a yellow sign with a black hazard symbol”, removing any loaded interpretation.
Appendix F Beyond Visual Recognition
While our primary focus was on classification tasks for the sake of clarity and control, we also extend our attacks to the image retrieval task using the Flickr30K dataset [24]. On Top-1 Image-to-Text retrieval, our attacks achieve an 83.4% success rate on Graphics w/ text, 63.8% success rate on Graphics wo/ text, and 63.4% on Text, demonstrating their ability to generalize to applications beyond classification.